Ty is a real estate agent and receives 2.5% commission on the first $200000, 1.5% on the next $100000 and 1% on the value thereafter. Find his commission for selling a property worth $825000


1
Expert's answer
2020-10-12T17:27:36-0400

"\frac{2.5}{100}\ \ of\ 200000\ = 5000"


"\frac{1.5}{100}\ \ of\ 100000 = 1500"


"remainder\ =\ 825000\ -\ (200000\ +\ 100000)"


"\frac{1}{100}\ \ of\ 525000 = 5250"


"5,000\ \ +\ 1500\ +\ 5250 = 11,750"


Ty will get a commission of $"11,\ 750"